The bob of a simple pendulum is made of wood. What will be the effect on the time
period if the wooden bob is replaced by an identical bob of iron?

Grade:7

2 Answers

Tanishq Malu
4 years ago
The time period of a simple pendulum is given by
  T=
                 2\pi \sqrt{\frac{l}{g}}
It is independent of any intrinsic property of a material. Thus, time period will remain same both for a wooden bob and an iron bob.
                                                                                                       
 
 
 
Arun
4 years ago
Since the sizes of metal bob and wooden bob are equal, the position of centre of mass does not
change. Therefore the length of the pendulum remains same so, time period remains constant.